-10(6-2)
-10 x 6 + -10 x -2
-60 x 20 =-120
-60 + 20 = 40
Not one of my answers on the multiple choice question. Thank you

Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

-10(6-2) = -10 x 6 + -10 x -2 = -60 + 20 = -40

Let me know if this helps.

Answer by solver91311(24713)








John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism